In thermodynamic process the pressure of a fixed mass of of a gas is changed such that the gas releases 20 j  of heat and 8 J work is done on the gas. If the initial internal energy of the gas was 30 J then the final internal and will be

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student
\Delta Q = \partial U+ \Delta W
-20 = (Uf - Ui) + (+8)
-28 = U- 30
U= 2 J
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT